OpenStudy (anonymous):

Add: 21.4 + (- 23.9)

Vocaloid (vocaloid):

hint: 21.4 + (- 23.9) = 21.4 - 23.9 = ?

Vocaloid (vocaloid):

(when you have something like A + (-B), it's the same as A - B)

OpenStudy (anonymous):

-2.5

Vocaloid (vocaloid):

great ~

OpenStudy (anonymous):

thank you